# Conceptualization of this study, Methodology, Software

### v2

@@ 2.2.4 Decoder
6:        for $i=1$ to $L$
7:              Compute IntraDDFE $\hat{H}^{\ell}_{s},\hat{H}^{\ell}_{a}$ with Equation (10).
8:              Compute InterDDFE $\tilde{H}^{\ell}_{s},\tilde{H}^{\ell}_{a}$ with Equation (11).
9:              Fuse the same feature and the output of the previous layer yields $H^{\ell}_{s}$ and $H^{\ell}_{a}$ with Equation (12) (13)
@@
2.3 Optimization
Predicting the drug-disease association can be viewed as a binary classification task. The objective is to determine whether a given drug is associated with a specific disease. Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function. We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively. We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($r_{i}$)-disease($d_{j}$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
@@ 2.3 Optimization
We optimize the DFDRNN model using the Adam (Kingma and Ba, 2014) optimizer and initialize the parameters’ weights with Xavier initialization (Glorot and Bengio, 2010). We randomly drop some edges and apply regular dropout (Srivastava et al., 2014) to improve the model’s generalization ability.
3 Experiments and Results
3.1 Baseline methods
To assess DFDRNN’s performance, we compared it with six state-of-the-art drug-disease association prediction models across the above four datasets. These models are:
$\bullet$ LAGCN (Yu et al., 2021): This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN (Cai et al., 2021): This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF (Meng et al., 2022): This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA (Zhang et al., 2024): This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT (Huang et al., 2024): This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN (Tang et al., 2023): This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
3.2 Experimental parameter configuration
@@
3.3 Evaluation metrics
We categorize the prediction results into four sets. As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative. We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
Table 2: The statistics of the four sets used in this study
Actual Predicted Positive Negative
Positive TP FN
Negative FP TN
[…]
$\text{FPR}=\frac{FP}{FP+TN}$ (20)
$\text{Precision}=\frac{TP}{TP+FP}$ (21)
Calculating the FPR, TPR, and precision at various thresholds, we construct the receiver operating characteristic (ROC) curve, with FPR on the X-axis and TPR on the Y-axis. Additionally, we generate the precision-recall (PR) curve, with recall on the X-axis and precision on the Y-axis. The evaluation metrics used for DFDRNN are the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
3.4 Performance of DFDRNN in the cross-validation
We employ the 10-fold cross-validation method to assess the model’s performance. During this process, the dataset is randomly partitioned into ten mutually exclusive subsets, each containing 10% known drug-disease association pairs and 10% unknown pairs. One subset serves as the test set, while the remaining nine subsets are used to train the model. This process is repeated ten times (i.e., 10-fold cross-validation) with different random seeds, and the average and standard deviation of the results are calculated to compare the model’s performance. The experimental results, including AUROC and AUPR values, are presented in Table 3. The ROC and PR curves can also be found in Supplementary Figures S1-S4.
In Table 3, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.75% and 3.1% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
Table 3: The AUROCs and AUPRs of DFDRNN and the other comparative models under 10-times 10-fold cross-validation.
Metrics Model Fdataset Cdataset Ldataset LRSSL Avg.
@@ 3.4 Performance of DFDRNN in the cross-validation
  • The best results in each column are in bold faces and the second best results are underlined.
3.5 Effect of the number of nearest neighbors
@@
In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during IntraDDFE, which influences the model’s performance. A small value of $t$ limits the neighbor information obtained, thereby affecting the model’s performance. On the other hand, as $t$ increases, more neighbors are obtained, but this also introduces noise information, leading to the propagation of unnecessary data. The performance of DFDRNN under 10-fold cross-validation varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention mechanism’s aggregation capability. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across the four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
3.6 Ablation analysis
3.6.1 Encoder
The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF). We introduce two dual-feature extraction modules, IntraDDFE and InterDDFE, to precisely encode drugs and diseases, with InterDDFE changing the feature (CF) of drugs (or diseases). We developed various model variants to underscore these elements’ importance in the DFDRNN encoder and evaluated their performance.
@@ 3.6.1 Encoder
  • DFDRNN-GCN: SAM was replaced by a GCN to observe the impact on the model’s performance.
  • DFDRNN-noCF: Remove the change of features in InterDDFE, meaning that the similarity feature and the association feature will not be changed into each other after InterDDFE.
Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
Variants Fdataset Cdataset Ldataset LRSSL
@@
Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-fold cross-validation. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Fdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Fdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the InterDDFE module.
3.6.2 Decoder
@@
The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Fdataset, Cdataset, Ldataset, and LRSSL. DFDRNN’s consistent outperformance in achieving the highest AUROC and AUPR values across all four datasets underscores the significance of the cross-dual-domain decoder. Furthermore, the performance of $s_{r}$, $s_{d}$, $s^{non}_{r}$, and $s^{non}_{d}$ consistently lags behind that of DFDRNN and $s^{non}$, further highlighting the effectiveness of dual-domain decoding.
@@
3.7 Identification of candidates for new diseases
@@
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic (Singh et al., 2020). Finding suitable drugs for emerging diseases is a complex and crucial task. To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset. Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples. The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
Figure 5: The performance of all methods in predicting potential drugs for new diseases on Fdataset: (a) Receiver operating characteristic (ROC) curves of prediction results obtained by applying DFDRNN and other competitive methods; (b) precision-recall (PR) curves of prediction results obtained by applying DFDRNN and other competitive methods.
@@
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.99%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
3.8 Case study
@@
To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset. We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs. Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD. To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB (Knox et al., 2024), PubChem (Kim et al., 2016), CTD (Davis et al., 2023), DrugCentral (Avram et al., 2021), and ClinicalTrials.
@@
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment (Erkulwater and Pillai, 1989). Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients (Lemke, 1995). DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
@@
Table 5: The top 10 DFDRNN-predicted candidate drugs for AD
@@ 10 DB01262 Decitabine CTD
Table 6: The top 10 DFDRNN-predicted candidate drugs for PD
@@ 10 DB00215 Citalopram CTD
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
4 Conclusion
Our study introduces DFDRNN, a neural network model designed for drug repositioning. DFDRNN uses IntraDDFE and InterDDFE modules to extract the similarity and association features of drugs (or diseases), respectively, achieving precise encoding of drugs (or diseases). Compared with the IntraDDFE module, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. In both modules, we use the self-attention mechanism to aggregate the neighbor information of drugs (and diseases). Additionally, the DFDRNN model fuses three similarity (or association) features into a new similarity (or association) feature and uses a layer attention mechanism to calculate the final similarity (or association) feature. Finally, a cross-dual-domain decoder process is adopted to predict drug-disease associations in both the drug and disease domains. DFDRNN performs best in 10-fold cross-validation across four benchmark datasets and achieves the highest AUROC in leave-one-out cross-validation. The experimental results confirm the effectiveness of our proposed DFDRNN model in drug repositioning.
For future work, we would like to investigate whether specific enhancements, such as integrating temporal data and handling multi-modal data, could further enhance the performance of DFDRNN. By incorporating temporal data, the model can better understand and predict patterns over time. Integration of data from different modalities could provide a more comprehensive understanding of the input. We aim to identify and implement the appropriate techniques and models for these enhancements. Furthermore, we intend to explore the integration of additional biological entities, such as miRNA and proteins, for drug repositioning.

### v3

@@ II-B4 Decoder
6:        for $\ell=$ to $L-1$
7:              Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
8:              Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
9:              Fuse the same feature and the input feature yields $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$ with Equation (12) (13)
@@
II-C Optimization
Predicting drug-disease associations can be treated as a binary classification problem, where the goal is to determine if a specific drug is associated with a specific disease. Due to the significant imbalance between the number of positive samples (associations) and negative samples (non-associations), we utilize binary cross-entropy as the loss function. Let $y^{+}$ and $y^{-}$ represent the positive and negative labels in the dataset, respectively. To mitigate the effects of this imbalance, we introduce a weight $\lambda=\frac{|y^{-}|}{|y^{+}|}$, where $|y^{-}|$ and $|y^{+}|$ denote the sizes of the negative and positive sample sets. A drug-disease pair $(i,j)$ corresponds to $r_{i}$ (drug) and $d_{j}$ (disease) in either $y^{+}$ or $y^{-}$, respectively.
@@ II-C Optimization
The DFDRNN model is optimized using the Adam optimizer [35], with parameters initialized through Xavier initialization [36]. To enhance generalization, we randomly drop edges and apply regular dropout techniques [37].
III Experiments and Results
III-A Baseline methods
We compared DFDRNN against six state-of-the-art methods for predicting drug-disease associations across the above four datasets. These methods include:
$\bullet$ LAGCN [17]: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN [18]: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF [38]: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA [21]: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT [23]: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN [20]: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
III-B Experimental parameter configuration
@@
TABLE II: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
Metrics Model Gdataset Cdataset Ldataset LRSSL Avg.
@@
  • The highest values in each column are highlighted in bold, while the second-highest values are underlined
III-C Evaluation metrics
We classify the prediction outcomes into four categories. As illustrated in Table III, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative. Based on these definitions, we proceed to calculate the True Positive Rate (TPR or Recall), False Positive Rate (FPR), and Precision using the following formulas:
TABLE III: The statistics of the four categories
Actual Predicted Positive Negative
Positive TP FN
Negative FP TN
[…]
$\text{FPR}=\frac{FP}{FP+TN}$ (20)
$\text{Precision}=\frac{TP}{TP+FP}$ (21)
By calculating the FPR, TPR, and precision across different thresholds, we generate the receiver operating characteristic (ROC) curve, plotting FPR on the X-axis and TPR on the Y-axis. Similarly, the precision-recall (PR) curve is created, with recall on the X-axis and precision on the Y-axis. The evaluation metrics for DFDRNN include the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
III-D DFDRNN’s performance in cross-validation
We utilize 10-fold cross-validation (10-CV) to evaluate the model’s performance. The dataset is randomly divided into ten distinct subsets, each comprising 10% known drug-disease associations and 10% unknown pairs. In each iteration, one subset is used as the test set, while the other nine serve as the training set. This process is repeated ten times with varying random seeds, allowing for the calculation of the average and standard deviation of the performance metrics. The experimental outcomes, including AUROC and AUPR scores, are summarized in Table II, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
In Table II, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.7% and 1.8% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
III-E Effect of the number of nearest neighbors
@@
In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during SDDFE, which influences the model’s performance. A small value of $t$ limits the neighbor information obtained, thereby affecting the model’s performance. On the other hand, as $t$ increases, more neighbors are obtained, but this also introduces noise information, leading to the propagation of unnecessary data. The performance of DFDRNN under 10-CV varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, demonstrating notable stability. This stability can be attributed to the mining capability of the SAM. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across the four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
III-F Ablation analysis
III-F1 Encoder
The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF). We introduce two dual-feature extraction modules, SDDFE and CDDFE, to precisely encode drugs and diseases, with CDDFE changing the feature (CF) of drugs (or diseases). We developed various model variants to underscore these elements’ importance in the DFDRNN encoder and evaluated their performance.
@@ III-F1 Encoder
  • DFDRNN-GCN: Replace the SAM with a GCN to evaluate its effect on the model’s performance.
  • DFDRNN-noCF: Remove the change of features in CDDFE, meaning that the similarity feature and the association feature will not be changed into each other after CDDFE.
Table IV presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Gdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Gdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the CDDFE module.
TABLE IV: The performance of DFDRNN and its four variants
Variants Gdataset Cdataset Ldataset LRSSL
@@
III-F2 Decoder
@@
The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Gdataset, Cdataset, Ldataset, and LRSSL. DFDRNN’s consistent outperformance in achieving the highest AUROC and AUPR values across all four datasets underscores the significance of the cross-dual-domain decoder. Furthermore, the performance of $s_{r}$, $s_{d}$, $s^{non}_{r}$, and $s^{non}_{d}$ consistently lags behind that of DFDRNN and $s^{non}$, further highlighting the effectiveness of dual-domain decoding.
@@
III-G Identification of candidates for new diseases
@@
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic [39]. Finding suitable drugs for emerging diseases is a complex and crucial task. To assess DFDRNN’s ability to predict candidate drugs for new diseases, we applied leave-one-out cross-validation (LOOCV) using the Gdataset. For each disease $d_{x}$, all drug-disease associations linked to $d_{x}$ were removed and treated as test data, while the remaining associations served as training data. The prediction relied on the similarity between $d_{x}$ and other diseases to identify its candidate drugs.
Figure 5: The performance of all approaches in predicting candidate drugs for new diseases on Gdataset: (a) ROC curves illustrating the prediction outcomes of DFDRNN alongside other competitive models; (b) PR curves illustrating the prediction outcomes of DFDRNN alongside other competitive models.
@@
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Gdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.81%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
III-H Case study
@@
To evaluate the practical effectiveness of DFDRNN, we performed case studies on two diseases, Alzheimer’s disease (AD) and Parkinson’s disease (PD), from the Gdataset. Using all drug-disease associations within the Gdataset as the training data, we trained DFDRNN to predict association probabilities for all unknown drug-disease pairs. We then ranked the drugs in descending order of predicted probabilities, identifying the top 10 candidate drugs for AD and PD. To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank [25], PubChem [40], CTD [28], DrugCentral [41], and ClinicalTrials.
@@
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table V displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment [42]. Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients [43]. DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table V have been confirmed by relevant authoritative data sources, with a success rate of 100%.
@@
TABLE V: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score
@@ 10 DB01262 Decitabine CTD
TABLE VI: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
@@ 10 DB00215 Citalopram CTD
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table VI presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
IV Conclusion
Our study introduces DFDRNN, a neural network model designed for drug repositioning. DFDRNN uses the SDDFE and CDDFE modules to extract the similarity and association features for drugs (or diseases), respectively, ensuring more appropriate encoding of drugs (or diseases). Compared with the SDDFE module, the CDDFE module changes the similarity (or association) feature into the association (or similarity) feature. In both modules, we use the SAM to mining the neighbor information of drugs (and diseases). Additionally, the DFDRNN model fuses three similarity (or association) features into a new similarity (or association) feature and uses a layer attention mechanism to calculate the final similarity (or association) feature. Finally, a cross-dual-domain decoder process is adopted to predict drug-disease associations in both the drug and disease domains. DFDRNN performs best in 10-CV across four benchmark datasets and achieves the highest AUROC in LOOCV. The experimental results confirm the effectiveness of our proposed DFDRNN in drug repositioning.
For future work, we would like to investigate whether specific enhancements, such as integrating temporal data and handling multi-modal data, could further enhance the performance of DFDRNN. By incorporating temporal data, the model can better understand and predict patterns over time. Integration of data from different modalities could provide a more comprehensive understanding of the input. We aim to identify and implement the appropriate techniques and models for these enhancements. Furthermore, we plan to investigate incorporating other biological entities, such as miRNAs and proteins, for drug repositioning.